1.  Below are the cash transactions of Kenya Co. for December 31, 2019:

          Cash payment to suppliers of goods and services              P 25,000

          Cash receipts from rendering of services                               50.000

          Payment to owners                                                      5,000

          Proceeds from bank loans                                             25,000

          Proceeds from cash investment of owners                      115,000

          Payment for acquisition of equipment                          100,000

          Cash at the beginning of the year                          1,000,000            

     

     Requirements:

     1.  How much is cash flow from operating activities?

     2.  How much is cash flow from investing activities?

     3.  How much is cash flow from financing activities?

     4.  How much is the net increase or decrease in cash for the year?

     5.  How much is cash as of December 31, 2019.

 



1
Expert's answer
2021-03-03T08:00:53-0500

1.) CASH FLOW FROM OPERATIONS

Cash received - Services. 50,000

Less cash paid - Suppliers. 25,000

Net cash flow from operations. 25,000


2.) CASH FLOW FROM INVESTING ACTIVITIES

Proceeds from investment. 115,000

Less. Purchase of equipment. 100,000

Net cash flow from investment 15,000


3.) CASH FLOW FROM FINANCING ACTIVITIES

Proceed from bank loan. 25,000

Less Payment to owners. 5,000

Net cash flow from financing. 20,000


4.) NET CASH INCREASE/DECREASE

Net cash increase/decrease= cash flow from operations+cash flow from investment+ cash flow from financing

Net cash increase= 25,000+15,000+20,000 = 60,000


5.) CASH AS AT DECEMBER 31,2019

Cash at end of year= Net increase in cash + Cash at beginning of year

Cash at year end= 60,000+1000 = 61,000